Answer in three decimal places


1.If block A has a velocity of 8.1 ft/sec to the right, determine the velocity of cylinder B. The velocity of B is positive if downward, negative if upward.


vB = __ ft/sec


2.At a certain instant, the velocity of cylinder B is 0.91 m/s down and its acceleration is 3.69 m/s2 up. Determine the corresponding velocity and acceleration of block A. The velocity and acceleration of A are positive if down the slope and negative if up the slope.


vA= _ m/s

aA= _ m/s2

1
Expert's answer
2021-04-22T07:27:19-0400

1)

"v_B=\frac{v_A}{2}=4.050~\frac{ft}{s}."

2)

"v_A=2v_B=1.820~\frac{m}{s},"

"a_A=\frac{a_B}{4}=0.923~\frac{m}{s^2}."